# Scalable Temporal Motif Densest Subnetwork Discovery

### v1

@@ Appendix A Application scenarios for the TMDS problem
(i) A travel network can be modeled as a temporal network $T=(V,E)$ with $V$ being the set of vertices corresponding to Points Of Interest (POIs) or particular geographic areas of a city (Lei et al., 2020), and edges of the form $(u,v,t)\in E$ represent trips by user that travel from point $u\in V$ to point $v\in V$ at time $t$. A temporal motif $M$ on such network captures therefore travel patterns and their dynamics. As an example a temporal motif $M:x\stackrel{{\scriptstyle t_{1}}}{{\to}}y\stackrel{{\scriptstyle t_{2}}}{{\to}}x$ occurring within one day often corresponds to a trip by a user from $x$ being its home, to $y$ being its work location, and clearly the user goes back home after work. The TMDS problem formulation, in such a scenario can provide unique insights on the POIs appearing frequently together in the various travel patterns (isomorphic to $M$) of the various users. Furthermore, analyzing different time-scales as captured by the temporal motif duration, can yield unique insights about daily vs. weekly vs. monthly patterns. As for our Figure 1 in the TMDS several POIs often coexist, and such information can be used to improve connections between POIs that are not well connected by public transport, or the for the design of travel passes for specific areas of a city with specific time-duration, based on POIs that are often visited together in the various trips.
(ii) E-commerce online platforms can be modeled as temporal bipartite networks $T=(V,E)$, where the set of vertices is partitioned in two layers $V=U\cup P$, with $U$ being the set of users (i.e., customers that purchase online products), and $P$ the set of products available for purchase. A temporal edge $e=(u,p,t)$ on such a network captures that user $u\in U$ purchased product $p\in P$ at a given time $t$. As an example a temporal motif $M^{\prime}:x\stackrel{{\scriptstyle t_{1}}}{{\to}}y,x\stackrel{{\scriptstyle t_{2}}}{{\to}}z$ occurring within a week corresponds to a pair of products i.e., $y$ and $z$ that user $x$ buys at distance of at most one week, note also that $z$ is purchased after $y$. A temporal motif $M$ therefore captures specific purchase habits of users within a limited time-limit (as again controlled by the duration parameter of the temporal motif). The optimal TMDS can be used to collect a set of users (and items) that frequently co-occur. In particular, the TMDS can contain several vertices (i.e., users) with similar purchase sequences or buying habits, this enables the design of personalized advertisement for those users in the TMDS (e.g., by leveraging the history of other users in the TMDS), furthermore this can be studied at different time-scales. Note also that this is more powerful than only considering products that are frequently co-purchased together (e.g., consider $M^{\prime}$, then the products $y$ and $z$ are purchased in different moments), for which many techniques already exist.

### v2

@@ Appendix A Application scenarios for the TMDS problem
(i) A travel network can be modeled as a temporal network $T=(V,E)$ with $V$ being the set of vertices corresponding to Points Of Interest (POIs) or particular geographic areas of a city (Lei et al., 2020), and edges of the form $(u,v,t)\in E$ represent trips by user that travel from point $u\in V$ to point $v\in V$ at time $t$. A temporal motif $M$ on such network captures therefore travel patterns and their dynamics. As an example a temporal motif $M:x\stackrel{{\scriptstyle t_{1}}}{{\to}}y\stackrel{{\scriptstyle t_{2}}}{{\to}}x$ occurring within one day often corresponds to a round trip by a user from $x$ being the home location, to $y$ being the work location. The TMDS problem formulation, in such a scenario can provide unique insights on the POIs appearing frequently together in the various travel patterns (isomorphic to $M$) of the various users. Furthermore, analyzing different time-scales as captured by the temporal motif duration, can yield unique insights about daily vs. weekly vs. monthly patterns. As for our Figure 1 in the TMDS several POIs often coexist, and such information can be used to improve connections between POIs that are not well connected by public transport, or for the design of travel passes for specific areas of a city with specific time-duration, based on POIs that are often visited together in the various trips.
(ii) E-commerce online platforms can be modeled as temporal bipartite networks $T=(V,E)$, where the set of vertices is partitioned in two layers $V=U\cup P$, with $U$ being the set of users (i.e., customers that purchase online products), and $P$ the set of products available for purchase. A temporal edge $e=(u,p,t)$ on such a network captures that user $u\in U$ purchased product $p\in P$ at a given time $t$. As an example a temporal motif $M^{\prime}:x\stackrel{{\scriptstyle t_{1}}}{{\to}}y,x\stackrel{{\scriptstyle t_{2}}}{{\to}}z$ occurring within a time-limit corresponds to a pair of products i.e., $y$ and $z$ that user $x$ buys at distance of at most the provided time-limit, note also that $z$ is purchased after $y$. A temporal motif $M$ therefore captures specific purchase habits of users within a limited time-limit (as again controlled by the duration parameter of the temporal motif). The optimal TMDS can be used to collect a set of users (and items) that frequently co-occur. In particular, the TMDS can contain several vertices (i.e., users) with similar purchase sequences or buying habits, this enables the design of personalized advertisement for those users in the TMDS (e.g., by leveraging the history of other users in the TMDS), furthermore this can be studied at different time-scales. Note also that this is more powerful than only considering products that are frequently co-purchased together (e.g., consider $M^{\prime}$, then the products $y$ and $z$ are purchased in different moments), for which many techniques already exist.